OpenStudy (anonymous):

what is the product in simplest form? State any restrictions on the variable. x^2+7x+10/x+3 * x^2-3x-18/x^2+x-2

OpenStudy (anonymous):

Factorise (too tired to use ^ so will use x2 for x^2) x2 + 7x + 10 = (x+5) (x+2) x2 - 3x - 18 = (x+3)(x-6) x2 + x - 2 = (x+2)(x-1) Put them in the equation, and cancel like terms. X CANT take any values which will make the denominator (bottom part) 0.

OpenStudy (anonymous):

i got that part already i just cant figure oout how to put them in equations.

OpenStudy (anonymous):

@phi help?

OpenStudy (phi):

NNL shows the final answer. notice that x ≠ -1, -2, or 1

OpenStudy (anonymous):

thank you!